# Ultralight Dark Matter Constraints from NanoHertz Gravitational Waves

### v3

@@ II.2 Orbital Decay Rate
where $\mu$, $\alpha_{\mu}$, and $\epsilon_{\mu}$ parameterise the relationship between the black hole mass ($M_{\mathrm{BH}}$) and the bulge mass ($M_{\mathrm{bulge}}$). Here, we set $\alpha_{\mu}$ to 1.1 following the fiducial value given in Table 2 of Agazie2023 while $\mu$ varies in our analysis. The Gaussian random scatter is denoted by $\mathcal{N}(0,\epsilon_{\mu})$; it has zero mean and its standard deviation $\epsilon_{\mu}$ is also varied, as discussed in Section IV. The fraction of the galaxy’s stellar mass present in the stellar bulge is $M_{\mathrm{bulge}}$, and it is related to the overall stellar mass of the galaxy $M_{\star}$ by
@@
$M_{c}=\frac{1}{4a^{1/2}}\left(\frac{\zeta(z)}{\zeta_{0}(z)}\right)^{1/6}\left(\frac{M_{\mathrm{ULDM}}}{M_{\mathrm{min},0}}\right)^{1/3},$ (6)
$\zeta(z)=\frac{18\pi^{2}+82(\Omega_{m}(z)-1)-39(\Omega_{m}(z)-1)^{2}}{\Omega_{m}(z)}$ (7)
@@
$M_{\mathrm{min},0}=\frac{32\pi\zeta(0)^{1/4}}{375^{1/4}(H_{0}m/\hbar)^{3/2}\Omega_{m0}^{3/4}}\rho_{m0}\,$, (8)
@@
In simulations of halo collapse in mixed models [44] with $m=2.5\times 10^{-22}$eV and halo masses of ${\cal{O}}(10^{9})M_{\odot}$ solitons do not form if $f\lesssim 0.3$. In the very largest halos (which are our primary interest), the central soliton will be particularly compact, and the lower bound on $f$ will presumably be smaller. We take $f\geq 0.01$ in what follows, which is a somewhat arbitrary boun,d but our results do not depend on its specific value. Note too that the definition of $f$ in Ref. [44] does not have the factor of ${\Omega_{\mathrm{DM}}}/{\Omega_{\mathrm{M}}}$ as baryons are not included in that analysis.
@@
where subscript $i$ indicates values associated with one of the black holes. The coefficient of friction $C_{i}$ is
@@
where subscript $j$ indicates the other black hole. This provides the torque, {align} dLidt=dridtdLidri
@@
If $e^{i\beta t}F(r)$ solves the radial Schrödinger-Poisson equations, the scaled profile $e^{i\gamma\beta t}\gamma F(\sqrt{\gamma}r)$ is also a solution. After restoring dimensionality, $\gamma$ is fixed by requiring the integrated mass (recalling that the density is proportional to $|\psi|^{2}$) matches the desired soliton mass, and the black hole to soliton mass ratio is invariant with $\gamma$. By varying the black hole mas,s we can solve iteratively for the profile [^1] In practice, we generated profiles for mass ratios between 0 and 82.15 and interpolated between them. and, as noted earlier, the pinching is significant when the black hole mass approaches or exceeds the soliton mass.
@@ V Results and Analysis
Figure 8: Comparisons of the soliton (orange) and NFW profile (blue) densities for a galaxy corresponding to $M_{\mathrm{BH}}=2.06\times 10^{8}\text{M}_{\odot}$ at z=0. Red vertical lines correspond to the frequencies of the PTA band, while the green vertical line is the core radius of the soliton.
@@ VI Discussion and Conclusion
We developed two models of the dynamical friction - the “realistic” scenario accounts for the SMBH binary pinching the soliton profile, whereas the “simple” one does not. The best-fit spectra for both models match the data with a similar level of quality. In both cases, scaling relations in the model mean that drag drops as a small, fractional power of the ULDM fraction. As a result, we cannot put a clear lower bound on the ULDM fraction. That said, the model relies on a number of hypotheses, and the soliton mass is set via the “standard” core-halo relation. Critically, this relation is an almost entirely unexplored relationship in models with a central SMBH, baryons and a mixture of ULDM and CDM. Each of these questions is an important avenue for future investigation. We have also assumed that there is a single ULDM component, which need not be the case, given that the arguments used to produce light scalar fields will typically be consistent with a spectrum with several such species [20].
Constraints on the stochastic background are expected to improve substantially as higher-quality PTA data becomes available in the next few years. If the data continues to suggest an absence of power at low frequencies (relative to the simple gravitational-wave–only SMBHB background prediction) PTA observations will identify a preferred region of the ULDM parameter space. Conversely, if this feature is not present in future datasets PTA observations will put stringent constraints on any possible ULDM contribution to the universe.
In the relativistic regime, the possibility of the binary-ULDM system forming a so-called ’gravitational atom’ for certain ULDM mass parameters has been considered [8], wherein excitation of the bound ULDM states produces an additional friction effect that may further impact the gravitational wave signal [22]. This additional effect may change our parameter constraints if added and is an interesting avenue for future work. Conversely, Ref. [14] considers the contribution of three-body scattering and eccentric orbits to the decay of SMBH binaries in the pulsar timing band and it will be critical to check that all environment effects associated with conventional components of the universe are fully accounted for.
@@ Acknowledgements.
We gratefully acknowledge support from the Marsden Fund Council grant MFP-UOA2131 from New Zealand Government funding, managed by the Royal Society Te Apārangi, and and we acknowledge the use of New Zealand eScience Infrastructure (NeSI) high-performance computing facilities. We thank Laura Burn, Emily Kendall, Renate Meyer, Pierre Mourier and Frank Wang for helpful discussions. We are grateful to Hovav Lazare for providing the code used to generate the other constraint contours in Fig. 4 and to the authors of holodeck for making the package available.

### v4

@@ II.2 Orbital Decay Rate
where $\mu$, $\alpha_{\mu}$, and $\epsilon_{\mu}$ parameterise the relationship between the black hole mass ($M_{\mathrm{BH}}$) and the bulge mass ($M_{\mathrm{bulge}}$). Here, we set $\alpha_{\mu}$ to 1.1 following the fiducial value given in Table 2 of Agazie2023 while $\mu$ varies in our analysis. The Gaussian random scatter is denoted by $\mathcal{N}(0,\epsilon_{\mu})$; it has zero mean and its standard deviation $\epsilon_{\mu}$ is also varied, as discussed in Section IV. The fraction of the galaxy’s stellar mass present in the stellar bulge is $M_{\mathrm{bulge}}$, and it is related to the overall stellar mass of the galaxy $M_{\star}$ by [36]
@@
$M_{c}=\frac{1}{4a^{1/2}}\left(\frac{\zeta(z)}{\zeta(0)}\right)^{1/6}\left(\frac{M_{\mathrm{ULDM}}}{M_{\mathrm{min},0}}\right)^{1/3}{M_{\mathrm{min},0}},$ (6)
$\zeta(z)=\frac{18\pi^{2}+82(\Omega_{\mathrm{M}}(z)-1)-39(\Omega_{\mathrm{M}}(z)-1)^{2}}{\Omega_{\mathrm{M}}(z)}$ (7)
@@
$M_{\mathrm{min},0}=\frac{32\pi\zeta(0)^{1/4}}{375^{1/4}(H_{0}m/\hbar)^{3/2}\Omega_{\mathrm{M0}}^{3/4}}\rho_{M0}\,$, (8)
@@
In simulations of halo collapse in mixed models [32] with $m=2.5\times 10^{-22}$eV and halo masses of ${\cal{O}}(10^{9})M_{\odot}$ solitons do not form if $f\lesssim 0.3$. In the very largest halos (which are our primary interest), the central soliton will be particularly compact, and the lower bound on $f$ will presumably be smaller. We take $f\geq 0.01$ in what follows, which is a somewhat arbitrary bound, but our results do not depend on its specific value. Note too that the definition of $f$ in Ref. [32] does not have the factor of ${\Omega_{\mathrm{DM}}}/{\Omega_{\mathrm{M}}}$ as baryons are not included in that analysis.
A separate consideration in mixed dark matter models is whether a soliton is able to reasonably form through the dynamical relaxation within the age of the universe. Following the formalism in [47], a rough criterion for this compares the radius in the initial halo that contains enough ULDM to form a soliton of the desired mass, $r_{supply}$, with the radius at which the relaxation timescale, given by [48]
$\tau=\frac{0.7\sqrt{2}m^{3}\sigma^{6}}{12\pi^{3}G^{2}\rho^{2}\textrm{ln}(m\sigma R)},$ (10)
is less than the age of the galaxy, $r_{relax}$. Here $\sigma$ is the velocity dispersion, $\rho$ is the initial ULDM density, and R is the characteristic radius of the system. Under the condition $r_{relax}>r_{supply}$, a soliton is expected to form.
This condition is not necessarily fulfilled for all of our parameter space - for black hole masses above $\sim 10^{8}\textrm{M}_{\odot}$ and particle masses $\sim 10^{-22}$eV, a majority of the dark matter must be ULDM for these conditions to be fulfilled. However, this treatment is for halos that have grown via direct collapse rather than mergers, as we consider here. Moreover, as noted above $f$ need not be directly equivalent to the ULDM fraction so we continue to treat it as a free parameter.
@@
where subscript $i$ indicates values associated with one of the black holes, and $\rho_{i}$ is the density of the ULDM soliton at the position of the black hole. The coefficient of friction $C_{i}$ is
@@
where subscript $j$ indicates the other black hole and $M_{\textrm{enc}_{i}}$ indicates the mass of soliton enclosed within the black hole’s position. This provides the torque, {align} dLidt=dridtdLidri
@@
If $e^{i\beta t}F(r)$ solves the radial Schrödinger-Poisson equations, the scaled profile $e^{i\gamma\beta t}\gamma F(\sqrt{\gamma}r)$ is also a solution. After restoring dimensionality, $\gamma$ is fixed by requiring the integrated mass (recalling that the density is proportional to $|\psi|^{2}$) matches the desired soliton mass, and the black hole to soliton mass ratio is invariant with $\gamma$. By varying the black hole mass, we can solve iteratively for the profile [^1] In practice, we generated profiles for mass ratios between 0 and 82.15 and interpolated between them. and, as noted earlier, the pinching is significant when the black hole mass approaches or exceeds the soliton mass.
@@
This model ignores the backreaction on the soliton structure that can be induced by the binary, which can cause oscillations in the orbital decay of the binary. However, with equal mass black holes the average decay is well-described by the semi-analytic formula [26]. That said, when one black hole is much larger than the other and both black hole masses are small compared to the soliton the excitation of dipole terms in the soliton can significantly slow the inspiral [49], but the systems that dominate our signal do not satisfy these conditions.
@@ V Results and Analysis
Figure 8: Comparisons of the soliton (orange) and NFW profile (blue) densities for a galaxy corresponding to $M_{\mathrm{BH}}=2.06\times 10^{8}\text{M}_{\odot}$ at z=0. Red vertical lines correspond to separations associated with gravitational wave emission in the PTA band, while the green vertical line is the core radius of the soliton.
@@ VI Discussion and Conclusion
We developed two models of the dynamical friction - the “realistic” scenario accounts for the SMBH binary pinching the soliton profile, whereas the “simple” one does not. The best-fit spectra for both models match the data with a similar level of quality. In both cases, scaling relations in the model mean that drag drops as a small, fractional power of the ULDM fraction. As a result, we cannot put a clear lower bound on the ULDM fraction. That said, the model relies on a number of hypotheses, and the soliton mass is set via the “standard” core-halo relation. Critically, this relation is an almost entirely unexplored relationship in models with a central SMBH, baryons and a mixture of ULDM and CDM. Each of these questions is an important avenue for future investigation. We have also assumed that there is a single ULDM component, which need not be the case, given that the arguments used to produce light scalar fields will typically be consistent with a spectrum with several such species [53]. We have also ignored the soliton coalescence timescale in fractional ULDM models [47]. In our case the halos are the products of mergers and $f$ itself is not directly equivalent to the ULDM fraction but this does suggest that there will be a lower bound on $f$ if additional drag is invoked to account for the PTA spectrum.
Constraints on the stochastic background are expected to improve substantially as higher-quality PTA data becomes available in the next few years. If the data continues to suggest an absence of power at low frequencies (relative to the simple gravitational-wave–only SMBH background prediction) PTA observations will identify a preferred region of the ULDM parameter space. Conversely, if this feature is not present in future datasets PTA observations will put stringent constraints on any possible ULDM contribution to the universe.
In the relativistic regime, the possibility of the binary-ULDM system forming a so-called “gravitational atom” for certain ULDM mass parameters has been considered [54], wherein excitation of the bound ULDM states produces an additional friction effect that may further impact the gravitational wave signal [55]. This additional effect may change our parameter constraints if added and is an interesting avenue for future work. Conversely, Ref. [56] considers the contribution of three-body scattering and eccentric orbits to the decay of SMBH binaries in the pulsar timing band and it will be critical to check that all environment effects associated with conventional components of the universe are fully accounted for.
@@ Acknowledgements.
We gratefully acknowledge support from the Marsden Fund Council grant MFP-UOA2131 from New Zealand Government funding, managed by the Royal Society Te Apārangi, and and we acknowledge the use of New Zealand eScience Infrastructure (NeSI) high-performance computing facilities. We thank Laura Burn, Emily Kendall, Renate Meyer, Pierre Mourier and Frank Wang for helpful discussions. We are grateful to Hovav Lazare for providing the code used to generate the other constraint contours in Fig. 4 and to the authors of holodeck for making the package available. We also acknowledge the University of Canterbury Research Cluster facilities for providing computational resources that significantly improved the efficiency of our computations (DOI:10.18124/CANTERBURYNZ-UCRCH, RRID:SCR_027870).
Data availibility
The code and the data used for producing the results of this paper are publicly avaliable [57].